# theorem]Theorem theorem]Corollary theorem]Lemma theorem]Example theorem]Assumption theorem]Proposition

### v4

@@ 4.2 Momentum based approach
SVRG in the non-convex setting. In particular, because Theorem 4.2 also applies for the case where we have multiple helpers, and we sample each time $S=1$ helpers, we get that SVRG converges in this case as $\mathcal{O}\big{(}\frac{(L+\delta K)F^{0}}{KT}\big{)}$, which matches the known SVRG rate (Reddi et al., 2016) (up to $\delta$ being small). More interestingly, we obtain the same convergence rate by using only one batch (no need to sample) if the batch is representative enough of the data (i.e., satisfies our Hessian similarity Assumption 3.3).
Local steps help in federated learning. By employing the decentralized version of this theorem (as detailed in Appendix C.4), we can ascertain that local steps (denoted as $K$ in our context) indeed provide a beneficial contribution to Federated Learning. This finding aligns with the results presented in (Karimireddy et al., 2020a).
@@ 4.3 MVR based approach
Convergence of this variant. We prove the following theorem that gives the convergence rate of this algorithm in the non-convex case. {framedtheorem} Under assumptions A3.1, 3.2,4.2. For ${\bm{m}}^{0}$ such that $E^{0}\leq\sigma_{f-h}^{2}/T$, for $a=\max(\frac{1}{T},1156\delta^{2}K^{2}\eta^{2})$ and $\eta=\min\Big{(}\frac{1}{L},\frac{1}{192\delta K},\frac{1}{K}\Big{(}\frac{F^{0}}{18432\delta^{2}T\sigma_{f-h}^{2}}\Big{)}^{1/3},\sqrt{\frac{F^{0}}{KT(L/2+8\delta K)}}\Big{)}\,.$ This choice gives us the rate :
@@
Baseline. Under the same assumptions and for the same amount of work, MVR or STORM (Cutkosky & Orabona, 2019) has the rate: $\mathcal{O}\bigg{(}\big{(}\frac{LF^{0}\sigma_{f}}{T}\big{)}^{2/3}+\frac{LF^{0}}{T}\bigg{)}$, this rate corresponds to needing at most $\mathcal{O}\bigg{(}\frac{LF^{0}\sigma_{f}}{\varepsilon^{3/2}}+\frac{LF^{0}}{\varepsilon}\bigg{)}$ (stochastic) gradient calls of $f$ to reach a point $\hat{{\bm{x}}}$ such $\mathbb{E}[\|\nabla f(\hat{{\bm{x}}})\|_{2}^{2}]\leq\varepsilon$ or $\varepsilon$-stationary point. This rate/iteration complexity is known to be optimal under the strong smoothness assumption: $f(\cdot,\xi)$ is $L$-smooth for all $\xi$ almost surely.
@@ 5 Potential applications
Reusing batches in SGD training and core-sets. In Machine Learning, the empirical risk minimization consisting of minimizing a function of the form $f({\bm{x}})=\frac{1}{N}\sum_{i=1}^{N}L({\bm{x}},\xi_{i})$ is ubiquitous. In many applications, we want to summarize the data-set $\{\xi_{i}\}_{i=1}^{N}$ by a smaller potentially weighted subset $CS=\{(\xi_{i_{j}},w_{j})\}_{j=1}^{M}$, for positive weights $(w_{j})_{j=1}^{M}$ that add up to one, this is referred to as a core-set (Bachem et al., 2017). In this case we can set $h({\bm{x}})=\sum_{(w,\xi)\in CS}wL({\bm{x}},\xi)$. An even sampler problem is when we sample a batch $B\subset\{\xi_{i}\}_{i=1}^{N}$ of size $b\leq N$, one question we can ask is how can we reuse this same batch to optimize $f$? In this case we set $h({\bm{x}})=(1/b)\sum_{\xi\in B}L({\bm{x}},\xi)$. In the case where we have many batches $\{B_{i}\}_{i\in I}$, we can set $h_{i}({\bm{x}})=(1/|B_{i}|)\sum_{\xi\in B_{i}}L({\bm{x}},\xi)$ for each $i\in I$ and use our decentralized framework to sample each time a helper $h$.
@@
Semi-supervised learning. In Semi-supervised Learning (Zhu, 2005), We have a small set of carefully cleaned data $\mathcal{Z}$ directly related to our target task and a large set of unlabeled data $\tilde{\mathcal{Z}}$. Let us also assume there is an auxiliary pre-training task defined over the source data, e.g., this can be the popular learning with contrastive loss (Chen et al., 2020). In this setting, we have a set of transformations $\mathcal{T}$ which preserves the semantics of the data, two unlabeled data samples $\tilde{\zeta_{1}},\tilde{\zeta_{2}}\in\tilde{\mathcal{Z}}$, and a feature extractor $\phi_{{\bm{x}}}(\cdot):\tilde{\mathcal{Z}}\rightarrow\mathbb{R}^{k}$ parameterized by ${\bm{x}}$. Then, the contrastive loss is of the form: ~$\tilde{\ell}(\,\phi_{{\bm{x}}}(\tilde{\zeta}_{1})\,,\,\phi_{{\bm{x}}}(\mathcal{T}(\tilde{\zeta_{1}}))\,,\,\phi_{{\bm{x}}}(\tilde{\zeta}_{2})\,)$ where the loss tries to minimize the distance between the representations $\phi_{{\bm{x}}}(\tilde{\zeta}_{1})$ and $\phi_{{\bm{x}}}(\mathcal{T}(\tilde{\zeta}_{1}))$, while simultaneously maximizing distance to $\phi_{{\bm{x}}}(\tilde{\zeta}_{2})$. Similarly, we also have a target loss $\ell:\mathcal{Z}\rightarrow\mathbb{R}$, which we care about; then, we can define
@@
Another way our framework can be used for semi-supervised learning is by endowing the unlabeled data with synthetic labels (either random labels as we will see for logistic regression or via Pseudo-labeling (Lee, 2013)) and defining the helper $h$ as the loss over the unlabeled data with its assigned labels.
Transfer learning. For a survey, see (Zhuang et al., 2020). In this case, in addition to a cleaned data set $\mathcal{Z}$ we have access to $\tilde{\mathcal{Z}}$ a pre-training source $\tilde{\mathcal{Z}}$. Given a fixed mask $M$ (for masking deep layers in a neural network) and a loss $\tilde{\ell}$, we set
$f({\bm{x}})=\mathbb{E}_{\zeta\in\mathcal{Z}}\big{[}{\ell({\bm{x}};\zeta)}\big{]}\quad\text{ and }\quad h({\bm{x}})=\mathbb{E}_{\tilde{\zeta}\in\tilde{\mathcal{Z}}}\big{[}\tilde{\ell}(M\odot{\bm{x}},\tilde{\zeta})\big{]}\,.$
Federated learning. Consider the problem of distributed/federated learning where data is decentralized over $N$ workers (McMahan et al., 2017b). Let $\{F_{1}({\bm{x}}),\dots,F_{N}({\bm{x}})\}$ represent the $N$ loss functions defined over their respective local datasets. In such settings, communication is much more expensive (say $M$ times more expensive) than the minibatch gradient computation time (Karimireddy et al., 2018). In this case, we set
@@
Thus, we want to minimize the target function $f({\bm{x}})$ defined over all the workers’ data but only have access to the cheap loss functions defined over local data. The main goal, in this case, is to limit the number of communications and use as many local steps as possible. Our proposed two variants are very close to MiMeSGDm and MiMeMVR from (Karimireddy et al., 2020a).
Personalized learning. This problem is a combination of the federated learning and transfer learning problems described above and is closely related to multi-task learning (Ruder, 2017). Here, there are $N$ workers, each with a task $\{F_{1}({\bm{x}}),\dots,F_{N}({\bm{x}})\}$, and without loss of generality, we describe the problem from the perspective of worker 1. In contrast to the delayed communication setting above, in this scenario, we only care about the local loss $F_{1}({\bm{x}})$, whereas all the other worker training losses $\{F_{2}({\bm{x}}),\dots,F_{N}({\bm{x}})\}$ constitute auxiliary data:
@@
Training with compressed models. Here, we want to train a large model parameterized by ${\bm{x}}\in\mathbb{R}^{d}$. To decrease the cost (both time and memory) of computing the backprop, we instead mask (delete) a large part of the parameters and perform backprop only on the remaining small subset of parameters (Sun et al., 2017; Yu & Huang, 2019). Suppose that our loss function $\ell$ is defined over sampled minibatches $\xi$ and parameters ${\bm{x}}$. Also, let us suppose we have access to some sparse/low-rank masks $\{\bm{1}_{\mathcal{M}_{1}},\dots,\bm{1}_{\mathcal{M}_{k}}\}$ from which we can choose. Then, we can define the problem as
@@ 6.1 Toy example
Figure 1: Effect of the bias $\zeta$ (zeta in the figure) on the naive approach (Naive), AuxMOM and Fine Tuning (FT) for $K=10$,$\delta=1$ and $\eta=\min(1/2,1/(\delta K))$. We can see that the naive approach fails to converge for large bias values, whereas AuxMOM converges all the time, no matter the value of the bias. Fine Tuning converges much slower for small values of $\delta$, but beats AuxMOM for $\delta=10$.
@@
Figure 2: Effect of the similarity $\delta$ (delta in the figure) on both the naive approach (Naive), AuxMOM and Fine Tuning for $K=10$,$\zeta=10$ and $\eta=0.5/(1+\delta)$ for Naive and AuxMOM and $\eta=0.5/(1+\delta)$ then $eta=0.5$ for Fine Tuning. We can see that the naive approach fails to benefit from small values of $\delta$; AuxMOM does not suffer from the same problem, whereas Fine Tuning is slower than AuxMOM.
@@ 6.2 Leveraging noisy or mislabeled data
Rotated features. We consider a simple feed-forward neural network ($Linear(28*28,512)\rightarrow ReLU\rightarrow Linear(512,512)\rightarrow ReLU\rightarrow Linear(512,10)$) to classify the MNIST dataset (LeCun & Cortes, 2010) , which is the main task $f$. As a helper function $h$, we rotate MNIST images by a certain angle $\in\{0,45,90,180\}$; the rotation plays the role of heterogeneity; it is worth noting that this is not simple data augmentation as numbers “meanings” are not conserved under such a transformation. We used a 256 batch size for $f$ and a $64$ batch size for $h$; in our experiments changing the batch size of $h$ to $256$ or $512$ led to similar results. We plot the test accuracy that is obtained using both the naive approach and AuxMOM.
@@
Figure 3: Effect of $K$ ($K-1$ is the number of times we use the helper h) on the test accuracy of the main task (for an angle = 45). We can see that our approach, as our theory predicts, benefits from bigger values of $K$.
@@ 6.4 Semi-supervised logistic regression
We consider a semi-supervised logistic regression task on the “Mushrooms” dataset from the libsvmtools repository (Chang & Lin, 2011), which has $8124$ samples, each with $112$ features. We divide this dataset into three equal parts: one for training and, one for testing, and the third one is unlabeled. In this context, the helper task $h$ is constituted of the unlabeled data to which we assigned random labels. Figure 10 shows indeed AuxMOM accelerates convergence on the training set. More importantly, it also leads to a smaller loss on the test set, which suggests a generalization benefit coming from using the unlabeled set.
@@ 7 Discussion
Higher order strategies. We believe our work can be “easily” extended to higher order strategies by using proxies of $f-h$ based on higher order derivatives. We expect that, in general, if we use a given Taylor approximation, we will need to make assumptions about its error. For example, if we use a 2nd order Taylor approximation, we expect we will need to bound the difference of the third derivatives of $f$ and $h$; this will be similar to the analysis of the Newton algorithm with cubic regularization(Nesterov & Polyak, 2006).
Dealing with the noise of the snapshot. In this work, we proposed to deal with the noise of the snapshot gradient of $f-h$ using momentum, other approaches such as using batch sizes of varying sizes with training (typically a batch size that increases as convergence is near) are possible, such an approach was used in SCSG(Lihua et al., ).
@@ C.1 Naive approach
As a reminder, the naive approach uses one unbiased gradient of $f$ at the beginning of each cycle followed by $K-1$ unbiased gradients of $h$.
0:  ${\bm{x}}_{0}$,${\bm{m}}^{0}$, $\eta$, $T$, $K$
@@
The proof is based on biased SGD theory. From (Ajalloeian & Stich, 2020) we have that if we the gradients ${\bm{g}}({\bm{x}})$ that we are using are such that :
@@
Summing the above inequality from $t=1,k=0$ to $t=T,k=K-1$ and then dividing by $KT$ we get the statement of the theorem. ∎
@@ C.2.4 Change during each cycle
Using the identity $-2ab=(a-b)^{2}-a^{2}-b^{2}$, we have :
@@
For $\eta\leq\frac{1}{192\delta K}$ we have $\delta^{2}\eta-\frac{\delta}{K}\leq 0$, $\eta+96K\delta\eta^{2}\leq 2\eta$ and $-\frac{\eta}{2}+48K\delta\eta^{2}\leq-\frac{\eta}{4}$ which gives the lemma. ∎
@@ C.3.1 Algorithm description
     ${\bm{m}}^{t}=(1-a){\bm{m}}^{t-1}+a{\bm{g}}^{f-h}+(1-a)({\bm{g}}^{f-h}-{\bm{g}}^{f-h}_{prev})$update momentum
@@ C.3.3 Proof
It is easy to verify that for $a\geq 144K^{2}\delta^{2}\eta^{2}$, we have $274K^{2}\delta^{2}\eta^{2}\leq a/$ 2 which finishes the proof. ∎
@@
About the need for Assumption C.9. Assumption C.9 becomes strong for $\delta$ small which is not good, as this should be the easiest case. however, we would like to point out that we only need Assumption C.9 to deal with the averaging that we perform to construct the new state ${\bm{x}}^{t}$. In the case where we sample each time one and only one helper (i.e. $S=1$), we don’t need such an assumption. This assumption was made in the context of Federated Learning for example in (Karimireddy et al., 2020a).

### v5

@@ 4.2 Momentum based approach
SVRG in the non-convex setting. In particular, because Theorem 4.2 also applies for the case where we have multiple helpers, and we sample each time $S=1$ helpers, we get that SVRG converges in this case as $\mathcal{O}\big(\frac{(L+\delta K)F^{0}}{KT}\big)$, which matches the known SVRG rate (SVRGnonC) (up to $\delta$ being small). More interestingly, we obtain the same convergence rate by using only one batch (no need to sample) if the batch is representative enough of the data (i.e., satisfies our Hessian similarity Assumption 3.3).
Local steps help in federated learning. By employing the decentralized version of this theorem (as detailed in Appendix C.4), we can ascertain that local steps (denoted as $K$ in our context) indeed provide a beneficial contribution to Federated Learning. This finding aligns with the results presented in (MiMe).
@@ 4.3 MVR based approach
Convergence of this variant. We prove the following theorem that gives the convergence rate of this algorithm in the non-convex case.
{framedtheorem}
Under assumptions A3.1, 3.2,4.2. For ${\bm{m}}^{0}$ such that $E^{0}\leq\sigma_{f-h}^{2}/T$, for $a=\max(\frac{1}{T},1156\delta^{2}K^{2}\eta^{2})$ and $\eta=\min\Big(\frac{1}{L},\frac{1}{192\delta K},\frac{1}{K}\Big(\frac{F^{0}}{18432\delta^{2}T\sigma_{f-h}^{2}}\Big)^{1/3},\sqrt{\frac{F^{0}}{KT(L/2+8\delta K)}}\Big)\,.$ This choice gives us the rate :
@@
Baseline. Under the same assumptions and for the same amount of work, MVR or STORM (MVR) has the rate: $\mathcal{O}\bigg(\big(\frac{LF^{0}\sigma_{f}}{T}\big)^{2/3}+\frac{LF^{0}}{T}\bigg)$, this rate corresponds to needing at most $\mathcal{O}\bigg(\frac{LF^{0}\sigma_{f}}{\varepsilon^{3/2}}+\frac{LF^{0}}{\varepsilon}\bigg)$ (stochastic) gradient calls of $f$ to reach a point $\hat{{\bm{x}}}$ such $\mathbb{E}[\|\nabla f(\hat{{\bm{x}}})\|_{2}^{2}]\leq\varepsilon$ or $\varepsilon$-stationary point. This rate/iteration complexity is known to be optimal under the strong smoothness assumption: $f(\cdot,\xi)$ is $L$-smooth for all $\xi$ almost surely.
@@ 5 Potential applications
Reusing batches in SGD training and core-sets. In Machine Learning, the empirical risk minimization consisting of minimizing a function of the form $f({\bm{x}})=\frac{1}{N}\sum_{i=1}^{N}L({\bm{x}},\xi_{i})$ is ubiquitous. In many applications, we want to summarize the data-set $\{\xi_{i}\}_{i=1}^{N}$ by a smaller potentially weighted subset $CS=\{(\xi_{i_{j}},w_{j})\}_{j=1}^{M}$, for positive weights $(w_{j})_{j=1}^{M}$ that add up to one, this is referred to as a core-set (coreset). In this case we can set $h({\bm{x}})=\sum_{(w,\xi)\in CS}wL({\bm{x}},\xi)$. An even sampler problem is when we sample a batch $B\subset\{\xi_{i}\}_{i=1}^{N}$ of size $b\leq N$, one question we can ask is how can we reuse this same batch to optimize $f$? In this case we set $h({\bm{x}})=(1/b)\sum_{\xi\in B}L({\bm{x}},\xi)$. In the case where we have many batches $\{B_{i}\}_{i\in I}$, we can set $h_{i}({\bm{x}})=(1/|B_{i}|)\sum_{\xi\in B_{i}}L({\bm{x}},\xi)$ for each $i\in I$ and use our decentralized framework to sample each time a helper $h$.
@@
Semi-supervised learning. In Semi-supervised Learning (zhu2005semi), We have a small set of carefully cleaned data $\mathcal{Z}$ directly related to our target task and a large set of unlabeled data $\tilde{\mathcal{Z}}$. Let us also assume there is an auxiliary pre-training task defined over the source data, e.g., this can be the popular learning with contrastive loss (chen2020simple). In this setting, we have a set of transformations $\mathcal{T}$ which preserves the semantics of the data, two unlabeled data samples $\tilde{\zeta_{1}},\tilde{\zeta_{2}}\in\tilde{\mathcal{Z}}$, and a feature extractor $\phi_{{\bm{x}}}(\cdot):\tilde{\mathcal{Z}}\rightarrow\mathbb{R}^{k}$ parameterized by $\bm{x}$}. Then, the contrastive loss is of the form: $\tilde{\ell}(\,\phi_{{\bm{x}}}(\tilde{\zeta}_{1})\,,\,\phi_{{\bm{x}}}(\mathcal{T}(\tilde{\zeta_{1}}))\,,\,\phi_{{\bm{x}}}(\tilde{\zeta}_{2})\,)$ where the loss tries to minimize the distance between the representations $\phi_{{\bm{x}}}(\tilde{\zeta}_{1})$ and $\phi_{{\bm{x}}}(\mathcal{T}(\tilde{\zeta}_{1}))$, while simultaneously maximizing distance to $\phi_{{\bm{x}}}(\tilde{\zeta}_{2})$. Similarly, we also have a target loss $\ell:\mathcal{Z}\rightarrow\mathbb{R}$, which we care about; then, we can define
@@
Another way our framework can be used for semi-supervised learning is by endowing the unlabeled data with synthetic labels (either random labels as we will see for logistic regression or via Pseudo-labeling (Pseudolabeling)) and defining the helper $h$ as the loss over the unlabeled data with its assigned labels.
Transfer learning. For a survey, see (TLsurvey). In this case, in addition to a cleaned data set $\mathcal{Z}$ we have access to $\tilde{\mathcal{Z}}$ a pre-training source $\tilde{\mathcal{Z}}$. Given a fixed mask $M$ (for masking deep layers in a neural network) and a loss $\tilde{\ell}$, we set
[…]
Federated learning. Consider the problem of distributed/federated learning where data is decentralized over $N$ workers (mcmahan2017communication). Let $\{F_{1}({\bm{x}}),\dots,F_{N}({\bm{x}})\}$ represent the $N$ loss functions defined over their respective local datasets. In such settings, communication is much more expensive (say $M$ times more expensive) than the minibatch gradient computation time (karimireddy2018adaptive). In this case, we set
@@
Thus, we want to minimize the target function $f({\bm{x}})$ defined over all the workers’ data but only have access to the cheap loss functions defined over local data. The main goal, in this case, is to limit the number of communications and use as many local steps as possible. Our proposed two variants are very close to MiMeSGDm and MiMeMVR from (MiMe).
Personalized learning. This problem is a combination of the federated learning and transfer learning problems described above and is closely related to multi-task learning (ruder2017overview). Here, there are $N$ workers, each with a task $\{F_{1}({\bm{x}}),\dots,F_{N}({\bm{x}})\}$, and without loss of generality, we describe the problem from the perspective of worker 1. In contrast to the delayed communication setting above, in this scenario, we only care about the local loss $F_{1}({\bm{x}})$, whereas all the other worker training losses $\{F_{2}({\bm{x}}),\dots,F_{N}({\bm{x}})\}$ constitute auxiliary data:
@@
Training with compressed models. Here, we want to train a large model parameterized by ${\bm{x}}\in\mathbb{R}^{d}$. To decrease the cost (both time and memory) of computing the backprop, we instead mask (delete) a large part of the parameters and perform backprop only on the remaining small subset of parameters (sun2017meprop; yu2019universally). Suppose that our loss function $\ell$ is defined over sampled minibatches $\xi$ and parameters $\bm{x}$}. Also, let us suppose we have access to some sparse/low-rank masks $\{\bm{1}_{\mathcal{M}_{1}},\dots,\bm{1}_{\mathcal{M}_{k}}\}$ from which we can choose. Then, we can define the problem as
@@ 6.1 Toy example
[…]
@@
[…]
@@ 6.2 Leveraging noisy or mislabeled data
Rotated features. We consider a simple feed-forward neural network ($Linear(28*28,512)\rightarrow ReLU\rightarrow Linear(512,512)\rightarrow ReLU\rightarrow Linear(512,10)$) to classify the MNIST dataset (mnist) , which is the main task $f$. As a helper function $h$, we rotate MNIST images by a certain angle $\in\{0,45,90,180\}$; the rotation plays the role of heterogeneity; it is worth noting that this is not simple data augmentation as numbers “meanings” are not conserved under such a transformation. We used a 256 batch size for $f$ and a $64$ batch size for $h$; in our experiments changing the batch size of $h$ to $256$ or $512$ led to similar results. We plot the test accuracy that is obtained using both the naive approach and AuxMOM.
@@
Figure 3: Effect of $K$ (K1K-1 is the number of times we use the helper h) on the test accuracy of the main task (for an angle = 45). We can see that our approach, as our theory predicts, benefits from bigger values of $K$.
@@ 6.4 Semi-supervised logistic regression
We consider a semi-supervised logistic regression task on the “Mushrooms” dataset from the libsvmtools repository (CC01a), which has $8124$ samples, each with $112$ features. We divide this dataset into three equal parts: one for training and, one for testing, and the third one is unlabeled. In this context, the helper task $h$ is constituted of the unlabeled data to which we assigned random labels. Figure 10 shows indeed AuxMOM accelerates convergence on the training set. More importantly, it also leads to a smaller loss on the test set, which suggests a generalization benefit coming from using the unlabeled set.
@@ 7 Discussion
Higher order strategies. We believe our work can be “easily” extended to higher order strategies by using proxies of fhf-h based on higher order derivatives. We expect that, in general, if we use a given Taylor approximation, we will need to make assumptions about its error. For example, if we use a 2nd order Taylor approximation, we expect we will need to bound the difference of the third derivatives of $f$ and $h$; this will be similar to the analysis of the Newton algorithm with cubic regularization(NewtonCubicReg).
Dealing with the noise of the snapshot. In this work, we proposed to deal with the noise of the snapshot gradient of fhf-h using momentum, other approaches such as using batch sizes of varying sizes with training (typically a batch size that increases as convergence is near) are possible, such an approach was used in SCSG(SCSG).
@@ C.1 Naive approach
As a reminder, the naive approach uses one unbiased gradient of $f$ at the beginning of each cycle followed by K1K-1 unbiased gradients of $h$.
[…]
@@
The proof is based on biased SGD theory. From (BiasedSGD) we have that if we the gradients ${\bm{g}}({\bm{x}})$ that we are using are such that :
@@
Summing the above inequality from $t=1,k=0$ to t=T,k=K1t=T,k=K-1 and then dividing by $KT$ we get the statement of the theorem. ∎
@@ C.2.4 Change during each cycle
Using the identity $2ab=(a-b)^{2}-a^{2}-b^{2}$, we have :
@@
For $\eta\leq\frac{1}{192\delta K}$ we have $\delta^{2}\eta-\frac{\delta}{K}\leq 0$, $\eta+96K\delta\eta^{2}\leq 2\eta$ and $\frac{\eta}{2}+48K\delta\eta^{2}\leq-\frac{\eta}{4}$ which gives the lemma. ∎
@@ C.3.1 Algorithm description
  ${\bm{m}}^{t}=(1-a){\bm{m}}^{t-1}+a{\bm{g}}^{f-h}+(1-a)({\bm{g}}^{f-h}-{\bm{g}}^{f-h}_{prev})$ §update momentum
@@ C.3.3 Proof
It is easy to verify that for $a\geq 144K^{2}\delta^{2}\eta^{2}$, we have $74K^{2}\delta^{2}\eta^{2}\leq a/2$ which finishes the proof. ∎
@@
About the need for Assumption C.9. Assumption C.9 becomes strong for $\delta$ small which is not good, as this should be the easiest case. however, we would like to point out that we only need Assumption C.9 to deal with the averaging that we perform to construct the new state ${\bm{x}}^{t}$. In the case where we sample each time one and only one helper (i.e. $S=1$), we don’t need such an assumption. This assumption was made in the context of Federated Learning for example in (MiMe).